f(x)=xsquared-2x+3
g(x)=xsquared-2x+5
c=-2
compute (fog)(c)

Answer by stanbon(75887)

f(x)=xsquared-2x+3
g(x)=xsquared-2x+5
c=-2
compute (fog)(c)
------------------------
fog(-2) = f[(-2)^2-2(-2)+5] = f[13] = 13^2-2*13+3 = 169-26+3 = 146
=================
Cheers,
Stan H.